Required information Use the following information for the Quick Study below. The following information applies to the questiTABLE B.1 p 1/(1i Present Value of 1 Rate 5% 8% 9% Perlods 1% 2% 3% 4% 6% 7% 10% 12% 15% 0.9434 1 0.9901 0.9804 0.9709 0.9615TABLE B.2 f= (1 Future Value of 1 Rate Perlods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% 1.0000 1.0000 1.0000 1.0000 1.0000 1.0TABLE B.3 /i (1i) Present Value of an Annuity of 1 Rate Perlods 1% 2% 3% 4% 5% 6% 7% 8% 9% 10% 12% 15% 0.9615 0.9901 0.9804TABLE B.4 f [(1 1Vi Future Value of an Annuity of 1 Rate 3% 12% Perlods 1% 2% 4% 5% 6% 7% 8% 9% 10% 15% 1.0000 1.0000 1.0000
Answer #1

Solution:

Annual depreciation = ($53,700 - $10,500) / 3 = $14,400

Annual cash inflows= Net income + Depreciation = $2,500 + $14,400 = $16,900

Cash Flow Select Chart Amount * PV factor = Present Value
Annual cash flow PVA of $1 $16,900.00 * 2.4869 = $42,029
Residual value PV of $1 $10,500.00 * 0.7513 = $7,889
Present value of cash inflows $49,917
Initial Investment $53,700
Net Present Value -$3,783